# Union of Finitely Generated Congruences on Ground Term Algebra

### v1

@@ 6 The Auxiliary dpwpa for the GTESs $E$ and $F$
By Definition 6.62, Statement 4.45, we have $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq n$. ∎
Intuitively if $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$, then there exists a $1$-context $\zeta$ over $\Sigma$ such that such that
@@
  • $R\langle E\cup F,\emptyset\rangle$ reach $a$ from $\zeta[b]$, and
  • $R\langle E\cup F,\emptyset\rangle$ reduces the subterms of $\zeta[b]$ at different prefixes of $\mathrm{addr}(\zeta)$ – except possibly for the pair of $\varepsilon$ and $\mathrm{addr}(\zeta)$ – to different constants in $C\langle E\cup F,\emptyset\rangle$.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Let $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$. Then there exists a $1$-context $\zeta$ over $\Sigma$, where
  • (a)
@@
[…]
@@
    $\zeta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$,
@@
    $\zeta[b]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[b]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
@@
Intuitively, the basic idea of the proof is similar to that of the pumping lemma for regular languages [18], context-free languages [18], and regular tree languages [12], [19]. The main difference from the proof of the pumping lemma is that we iterate contracting the ground term rather than pumping it. We stop when we get such a small ground term that we cannot contract it any more.
It is sufficient to prove that there exists a $1$-context $\zeta$ over $\Sigma$ such that Conditions (a)–(e) hold. Because for any $1$-context $\zeta$ over $\Sigma$ satisfying Conditions (a)–(e), Condition (f) holds as well. By Statement 4.43, there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$, i.e. Conditions (a)–(d) hold for $\zeta=\delta$. We now distinguish two cases:
@@
[…]
$\left(0\leq i<j\leq k-1\mbox{ or }1\leq i<j\leq k\right)$ and $\zeta[b]|_{\beta_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=\zeta[b]|_{\beta_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
@@
$\delta^{\prime}\neq\diamondsuit$, $\mathrm{length}(\mathrm{addr}(\delta^{\prime}))<\mathrm{length}(\mathrm{addr}(\delta))$, and $\delta^{\prime}[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$ i.e. Conditions (a)–(d) hold for $\zeta=\delta^{\prime}$.
@@
Intuitively, we now consider two constants, $a$ and $b$ in $C\langle E\cup F,\emptyset\rangle$ and a sequence of rewrite rules in $R\langle{E\cup F},\emptyset\rangle$ of which consecutive applications leads us from $b$ to $a$. Let $a,b\in C\langle E\cup F,\emptyset\rangle$, $k\geq 1$, and for each $j=1,\ldots,k$, let
  • $\sigma_{j}(a_{j1},\ldots,a_{jm_{j}})\dot{\rightarrow}a_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, be a rewrite rule and
[…]
@@
  • $b=a_{1i_{1}}$,
  • for each $j=1,\ldots,k-1$, $a_{j}=a_{(j+1)i_{j+1}}$, and
  • $a=a_{k}$.
The above sequence of rewrite rules is connected in the sense that we can apply them one after another, the right-hand side of each rule appears on the left-hand side of the next rule. Furthermore, the consecutive applications of these rewrite rules leads us from the constant $b$ to the constant $a$ since $b$ appears on the left-hand side of the first rule, and $a$ is the right-hand side of the last rule. We can apply the above rewrite rules one after another in the following way:
$b$ appears in the left-hand side of the first rewrite rule,
we apply the first rewrite rule to its left-hand side, $\sigma_{1}(a_{11},\ldots,a_{1m_{1}})$, we get the right-hand side $a_{1}$, which appears on the left-hand side of the second rewrite rule,
we apply the second rewrite rule to its left-hand side, $\sigma_{2}(a_{21},\ldots,a_{2m_{2}})$, we get the right-hand side $a_{2}$, which appears on the left-hand side of the third rewrite rule,
we apply the third rewrite rule to its left-hand side, $\sigma_{3}(a_{31},\ldots,a_{3m_{3}})$, we get the right-hand side $a_{3}$, which appears on the left-hand side of the fourth rewrite rule, and son on.
Finally, the right-hand side $a_{k-1}$ of the last but one rewrite rule appears in the left-hand side of the last rewrite rule,
we apply the last rewrite rule to its left-hand side, $\sigma_{k}(a_{k1},\ldots,a_{km_{k}})$, we get its right-hand side $a_{k}=a$.
We now state formally and show the following intuitive statement: if $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$, then there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from the constant $b$ to the constant $a$. Moreover, $b$ and the intermediate constants between $b$ and $a$, appearing as the right-hand sides of the applied rewrite rules except for the last one, are pairwise different.
@@
[…]
  • $b=a_{1i_{1}}$,
  • for each $j=1,\ldots,k-1$, $a_{j}=a_{(j+1)i_{j+1}}$,
  • $b\not\in\{\,a_{1},\ldots,a_{k-1}\,\}$ and for any $11\leq i<j\leq k-$ 1, $a_{i}\neq a_{j}$, and
  • $a=a_{k}$.
@@
[…]
@@
    $\zeta[b]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=a$,
@@
    $\zeta[b]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[b]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
@@
As $\mathrm{addr}(\xi)=\alpha_{k}$, we have $\xi[b]|_{\alpha_{k}}=b$.
For each $i\in\{1,\ldots,k-1\,\}$, let $a_{i}=\zeta[b]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
Since $b=\zeta[b]|_{\alpha_{k}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$, by Condition (e) we have
$b\not\in\{\,a_{1},\ldots,a_{k-1}\,\}$ and for any $11\leq i<j\leq k-$ 1, $a_{i}\neq a_{j}$.
@@
$\zeta[b]=t_{0}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{1}\mathop{\rightarrow}_{[\alpha_{k-1},\sigma_{1}(a_{11},\ldots,a_{1m_{1}})\dot{\rightarrow}a_{1}]}t_{1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{2}\mathop{\rightarrow}_{[\alpha_{k-2},\sigma_{2}(a_{21},\ldots,a_{2m_{2}})\dot{\rightarrow}a_{2}]}t_{2}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}\newline \cdots\mathop{\rightarrow}_{[\alpha_{1},\sigma_{k-1}(a_{(k-1)1},\ldots,a_{(k-1)m_{1}})\dot{\rightarrow}a_{k-1}]}t_{k-1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{k}\mathop{\rightarrow}_{[\alpha_{0},\sigma_{k}(a_{k1},\ldots,a_{km_{k}})\dot{\rightarrow}a_{k}]}a_{k}=a$.
@@
  • for each $i\in\{1,\ldots,k\,\}$, $a_{i}=\zeta[b]|_{\alpha_{k-i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$,
[…]
  • for each $j\in\{1,\ldots,k\,\}$, $\sigma_{j}(a_{j1},\ldots,a_{jm_{j}})\dot{\rightarrow}a_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$,
[…]
  • $a=a_{k}$. ∎
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. If $R\langle E\cup F,\emptyset\rangle$ reach a constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant $b\in C\langle E\cup F,\emptyset\rangle$, then there exists a path of positive length or a cycle from the vertex $a$ to the vertex $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
@@
We have the lemma by Lemma 6.65, Statement 4.43, and by the definition of $A[E;F]$ in Definition 6.62. ∎
We now show the converse of Lemma 6.66. Intuitively, if there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from a constant $b\in C\langle E\cup F,\emptyset\rangle$ to a constant $a\in C\langle E\cup F,\emptyset\rangle$, then $R\langle E\cup F,\emptyset\rangle$ reach the constant $a$ from a proper extension of the constant $b$.
Lemma 6.67.
[…]
$\sigma(s_{1},\ldots,s_{i-1},a_{i},s_{i+1},\ldots,s_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma(a_{1},\ldots,a_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a$.
Proof.
By Definition 4.26, Definition 4.28, and Proposition 4.32 we have the lemma. ∎
Lemma 6.68.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Let $k\geq 1$, and $(a_{k},a_{k-1}),\,(a_{k-1},a_{k-2}),\,\ldots,\,(a_{1},a_{0})\in A[E;F]$. Then there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[a_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}a_{k}$.
Proof.
By Definition 6.62 and Definition 4.44, for each $j=1,\ldots,k$, there exists a rewrite rule $\sigma_{j}(a_{j1},\ldots,a_{jm_{j}})\dot{\rightarrow}a_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, such that
[…]
@@
$\delta\neq\diamondsuit$ and $\delta[a_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}a_{k}$.
@@
$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}(a_{(k+1)1},\ldots,a_{(k+1)i_{k+1}-1},a_{(k+1)i_{k+1}},a_{(k+1)i_{k+1}+1},\ldots,a_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a_{k+1}$.
@@
$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},\delta[a_{0}],s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\newline \sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},a_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}(a_{(k+1)1},\ldots,a_{(k+1)i_{k+1}-1},a_{(k+1)i_{k+1}},a_{(k+1)i_{k+1}+1},\ldots,a_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}a_{k+1}$. ∎
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. If there exists a path of positive length or a cycle from a vertex $a$ to a vertex $b$ in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E\cup F,\emptyset\rangle$ reach the constant $a\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of the constant $b\in C\langle E\cup F,\emptyset\rangle$.
Proof.
We have the lemma by Lemma 6.68. ∎

### v2

@@ 6 The Auxiliary dpwpa for the GTESs $E$ and $F$
By Definition 6.63, Statement 4.46, we have $|A[E;F]|=|BSTEP\langle E\cup F,\emptyset\rangle|\leq 2n$. ∎
Intuitively, the basic idea of the proof of the following lemma is similar to that of the pumping lemma for regular languages [20], context-free languages [20], and regular tree languages [18], [21]. The main difference from the proof of the pumping lemma is that we iterate contracting the ground term rather than pumping it. We stop when we get such a small ground term that we cannot contract it any more. If $R\langle E\cup F,\emptyset\rangle$ reaches a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$, then there exists a $1$-context $\zeta$ over $\Sigma$ such that such that
  • 1.
@@
  • 2.
    $R\langle E\cup F,\emptyset\rangle$ reaches b{b} from $\zeta[{c}]$, and
  • 3.
    $R\langle E\cup F,\emptyset\rangle$ reduces the subterms of $\zeta[{c}]$ at different prefixes of $\mathrm{addr}(\zeta)$ – except possibly for the pair of $\varepsilon$ and $\mathrm{addr}(\zeta)$ – to different constants in $C\langle E\cup F,\emptyset\rangle$.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Let $R\langle E\cup F,\emptyset\rangle$ reach a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$. Then there exists a $1$-context $\zeta$ over $\Sigma$, where
  • (b)
@@
    for every $i\in\{\,0,\ldots,k\,\}$, $\alpha_{i}$ denotes the prefix of $\mathrm{addr}(\zeta)$ of length $i$
@@
    $\zeta[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}={b}$,
@@
    $\zeta[{c}]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[{c}]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
@@
It is sufficient to prove that there exists a $1$-context $\zeta$ over $\Sigma$ such that Conditions (a)–(e) hold. Because for any $1$-context $\zeta$ over $\Sigma$ satisfying Conditions (a)–(e), Condition (f) holds as well. By Statement 4.44, there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}={b}$, i.e. Conditions (a)–(d) hold for $\zeta=\delta$. We now distinguish two cases:
@@
Case 2: Condition (e) does not hold for $\zeta=\delta$. For every $i\in\{\,0,\ldots,n\,\}$, let $\beta_{i}$ denote the prefix of $\mathrm{addr}(\delta)$ of length $i$. There exist $i,j\in\{\,0,\ldots,k\,\}$ such that
$\left(0\leq i<j\leq k-1\mbox{ or }1\leq i<j\leq k\right)$ and $\zeta[{c}]|_{\beta_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}=\zeta[{c}]|_{\beta_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
@@
$\delta^{\prime}\neq\diamondsuit$, $\mathrm{length}(\mathrm{addr}(\delta^{\prime}))<\mathrm{length}(\mathrm{addr}(\delta))$, and $\delta^{\prime}[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}={b}$ i.e. Conditions (a)–(d) hold for $\zeta=\delta^{\prime}$.
@@
Intuitively, we now consider two constants, b{b} and c{c} in $C\langle E\cup F,\emptyset\rangle$ and a sequence of rewrite rules in $R\langle{E\cup F},\emptyset\rangle$ of which consecutive applications leads us from c{c} to b{b}. Let ${b},{c}\in C\langle E\cup F,\emptyset\rangle$, $k\geq 1$, and for each $j=1,\ldots,k$, let
  • 1.
    $\sigma_{j}({b}_{j1},\ldots,{b}_{jm_{j}})\dot{\rightarrow}{b}_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, be a rewrite rule and
  • 2.
    $i_{j}\in\{\,1,\ldots,m_{j}\,\}$ be an integer
@@
  • 1.
    ${c}={b}_{1i_{1}}$,
  • 2.
    for each $j=1,\ldots,k-1$, ${b}_{j}={b}_{(j+1)i_{j+1}}$, and
  • 3.
    ${b}={b}_{k}$.
The above sequence of rewrite rules is connected in the sense that we can apply them one after another, the right-hand side of each rule appears on the left-hand side of the next rule. Furthermore, the consecutive applications of these rewrite rules leads us from the constant c{c} to the constant b{b} since c{c} appears on the left-hand side of the first rule, and b{b} is the right-hand side of the last rule. We can apply the above rewrite rules one after another in the following way:
c{c} appears in the left-hand side of the first rewrite rule,
we apply the first rewrite rule to its left-hand side, $\sigma_{1}({b}_{11},\ldots,{b}_{1m_{1}})$, we get the right-hand side ${b}_{1}$, which appears on the left-hand side of the second rewrite rule,
we apply the second rewrite rule to its left-hand side, $\sigma_{2}({b}_{21},\ldots,{b}_{2m_{2}})$, we get the right-hand side ${b}_{2}$, which appears on the left-hand side of the third rewrite rule, and son on.
Finally, the right-hand side ${b}_{k-1}$ of the last but one rewrite rule appears in the left-hand side of the last rewrite rule,
we apply the last rewrite rule to its left-hand side, $\sigma_{k}({b}_{k1},\ldots,{b}_{km_{k}})$, we get its right-hand side ${b}_{k}={b}$.
We now state formally and show the following intuitive statement: if $R\langle E\cup F,\emptyset\rangle$ reaches a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$, then there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from the constant c{c} to the constant b{b}. Moreover, c{c} and the intermediate constants between c{c} and b{b}, appearing as the right-hand sides of the applied rewrite rules except for the last one, are pairwise different.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Let $R\langle E\cup F,\emptyset\rangle$ reach a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$. Then there exists an integer $1\leq k\leq|C\langle E\cup F,\emptyset\rangle|$ and for each $j=1,\ldots,k$, there exists a rewrite rule $\sigma_{j}({b}_{j1},\ldots,{b}_{jm_{j}})\dot{\rightarrow}{b}_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, and an integer $i_{j}\in\{\,1,\ldots,m_{j}\,\}$ such that
  • 1.
    ${c}={b}_{1i_{1}}$,
  • 2.
    for each $j=1,\ldots,k-1$, ${b}_{j}={b}_{(j+1)i_{j+1}}$,
  • 3.
    ${c}\not\in\{\,{b}_{1},\ldots,{b}_{k-1}\,\}$ and for any $1\leq i<j\leq k-1$, ${b}_{i}\neq{b}_{j}$, and
  • 4.
    ${b}={b}_{k}$.
@@
    for every $i\in\{\,0,\ldots,k\,\}$, $\alpha_{i}$ denotes the prefix of $\mathrm{addr}(\zeta)$ of length $i$
@@
    $\zeta[{c}]{\downarrow}_{R\langle E\cup F,\emptyset\rangle}={b}$,
@@
    $\zeta[{c}]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}\neq\zeta[{c}]|_{\alpha_{j}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$ for any $0\leq i<j\leq k$ such that $1\leq i$ or $j\leq k-1$, and
@@
As $\mathrm{addr}(\xi)=\alpha_{k}$, we have $\xi[{c}]|_{\alpha_{k}}={c}$.
For each $i\in\{1,\ldots,k-1\,\}$, let ${b}_{i}=\zeta[{c}]|_{\alpha_{i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$.
Since ${c}=\zeta[{c}]|_{\alpha_{k}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$, by Condition (e) we have
${c}\not\in\{\,{b}_{1},\ldots,{b}_{k-1}\,\}$ and for any $1\leq i<j\leq k-1$, ${b}_{i}\neq{b}_{j}$.
@@
$\zeta[{c}]=t_{0}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{1}\mathop{\rightarrow}_{[\alpha_{k-1},\sigma_{1}({b}_{11},\ldots,{b}_{1m_{1}})\dot{\rightarrow}{b}_{1}]}t_{1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{2}\mathop{\rightarrow}_{[\alpha_{k-2},\sigma_{2}({b}_{21},\ldots,{b}_{2m_{2}})\dot{\rightarrow}{b}_{2}]}t_{2}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}\newline \cdots\mathop{\rightarrow}_{[\alpha_{1},\sigma_{k-1}({b}_{(k-1)1},\ldots,{b}_{(k-1)m_{1}})\dot{\rightarrow}{b}_{k-1}]}t_{k-1}\mathop{\rightarrow}_{R\langle{E\cup F},\emptyset\rangle}^{*}s_{k}\mathop{\rightarrow}_{[\alpha_{0},\sigma_{k}({b}_{k1},\ldots,{b}_{km_{k}})\dot{\rightarrow}{b}_{k}]}{b}_{k}={b}$.
@@
  • 2.
    for each $i\in\{1,\ldots,k\,\}$, ${b}_{i}=\zeta[{c}]|_{\alpha_{k-i}}{\downarrow}_{R\langle E\cup F,\emptyset\rangle}$,
  • 3.
    ${c}={b}_{1i_{1}}$ for some $i_{1}\in\{\,1,\ldots,m_{1}\,\}$,
  • 4.
    for each $j\in\{1,\ldots,k\,\}$, $\sigma_{j}({b}_{j1},\ldots,{b}_{jm_{j}})\dot{\rightarrow}{b}_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$,
  • 5.
    for each $j=1,\ldots,k-1$, there exists $i_{j+1}\in\{\,1,\ldots,m_{j+1}\,\}$ such that ${b}_{j}={b}_{(j+1)i_{j+1}}$, and
  • 6.
    ${b}={b}_{k}$. ∎
We have the following result by Lemma 6.66, Statement 4.44, and by the definition of $A[E;F]$ in Definition 6.63.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. If $R\langle E\cup F,\emptyset\rangle$ reaches a constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of a constant ${c}\in C\langle E\cup F,\emptyset\rangle$, then there exists a path of positive length or a cycle from the vertex b{b} to the vertex c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$.
We now show the converse of Lemma 6.67. Intuitively, if there exists a connected sequence of rewrite rules in $R\langle E\cup F,\emptyset\rangle$ that leads us from a constant ${c}\in C\langle E\cup F,\emptyset\rangle$ to a constant ${b}\in C\langle E\cup F,\emptyset\rangle$, then $R\langle E\cup F,\emptyset\rangle$ reaches the constant b{b} from a proper extension of the constant c{c}. By Definition 4.27, Definition 4.29, and Proposition 4.33 we have the following lemma.
Lemma 6.68.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Let $\sigma({b}_{1},\ldots,{b}_{m})\dot{\rightarrow}{b}\in R\langle{E\cup F},\emptyset\rangle$ be a rewrite rule, where $m\geq 1$, ${b}_{1},\ldots,{b}_{m},{b}\in C\langle E\cup F,\emptyset\rangle$. Let $i\in\{\,1,\ldots,m\,\}$. Then there exist ground terms $s_{1},\ldots,s_{i-1},s_{i+1},\ldots,s_{m}\in T_{\Sigma}$ such that
$\sigma(s_{1},\ldots,s_{i-1},{b}_{i},s_{i+1},\ldots,s_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma({b}_{1},\ldots,{b}_{m})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{b}$.
Lemma 6.69.
Let $E$ and $F$ be GTESs over a signature $\Sigma$. Let $k\geq 1$, and $({b}_{k},{b}_{k-1}),\,({b}_{k-1},{b}_{k-2}),\,\ldots,\,({b}_{1},{b}_{0})\in A[E;F]$. Then there exists a $1$-context $\delta$ over $\Sigma$ such that $\delta\neq\diamondsuit$ and $\delta[{b}_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{b}_{k}$.
@@
By Definition 6.63 and Definition 4.45, for each $j=1,\ldots,k$, there exists a rewrite rule $\sigma_{j}({b}_{j1},\ldots,{b}_{jm_{j}})\dot{\rightarrow}{b}_{j}\in R\langle{E\cup F},\emptyset\rangle$, $m_{j}\geq 1$, such that
${b}_{0}\in\{\,{b}_{11},\ldots,{b}_{1m_{1}}\,\}$ and for each $j=1,\ldots,k-1$, there exists $i_{j+1}\in\{\,1,\ldots,m_{j+1}\,\}$ such that ${b}_{j}={b}_{(j+1)i_{j+1}}$.
@@
$\delta\neq\diamondsuit$ and $\delta[{b}_{0}]\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}{b}_{k}$.
@@
$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},{b}_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},{b}_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}({b}_{(k+1)1},\ldots,{b}_{(k+1)i_{k+1}-1},{b}_{(k+1)i_{k+1}},{b}_{(k+1)i_{k+1}+1},\ldots,{b}_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{b}_{k+1}$.
@@
$\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},\delta[{b}_{0}],s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},{b}_{k},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})=\newline \sigma_{k+1}(s_{1},\ldots,s_{i_{k+1}-1},{b}_{(k+1)i_{k+1}},s_{i_{k+1}+1},\ldots,s_{m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}^{*}\newline \sigma_{k+1}({b}_{(k+1)1},\ldots,{b}_{(k+1)i_{k+1}-1},{b}_{(k+1)i_{k+1}},{b}_{(k+1)i_{k+1}+1},\ldots,{b}_{(k+1)m_{k+1}})\mathop{\rightarrow}_{R\langle E\cup F,\emptyset\rangle}{b}_{k+1}$. ∎
We have the following statement by Lemma 6.69.
@@
Let $E$ and $F$ be GTESs over a signature $\Sigma$. If there exists a path of positive length or a cycle from a vertex b{b} to a vertex c{c} in the auxiliary dpwpa $AUX[E;F]=(C\langle E\cup F,\emptyset\rangle,A[E;F])$, then $R\langle E\cup F,\emptyset\rangle$ reaches the constant ${b}\in C\langle E\cup F,\emptyset\rangle$ from a proper extension of the constant ${c}\in C\langle E\cup F,\emptyset\rangle$.
By Lemma 6.67 and Lemma 6.70, we have the following result.